# RetrievalAttention: Accelerating Long-Context LLM Inference via Vector Retrieval

### v2

@@ A.4 Decoding Latency on A100
We test the generality of RetrievalAttention by measuring its performance on a server with one A100 (80GB) and one AMD EPYC CPU with 24 cores and 220GB DRAM. We show the token-generation latency of different methods on three models in Table 8. Since the KV cache of full attention is disabled, all prompt tokens need to be recalculated during the decoding, incurring a very high decoding latency. By enabling the KV cache with the PageAttention optimization in vLLM, the decoding latency is significantly reduced. However, vLLM suffers from OOM issue with the increase of context length, which we elaborate further later. Other KV cache dropping or block retrieval methods including StreamingLLM, SnapKV, and InfLLM achieve faster decoding speed, but this is at the expense of a significant drop in model accuracy. In contrast, RetrievalAttention does not compromise generation accuracy while achieving much lower decoding latency than IVF and Flat because of the efficient mitigation of out-of-distribution problem.
@@ B.1 Formula of Combining Attention Results from the CPU and GPU Side
Attention operation is applied to the two sets of KV vectors separately on CPU and GPU, generating two partial attention outputs (denoted as $\mathbf{o}_{\mathcal{W}}$ and $\mathbf{o}_{\Omega}$, respectively). To guarantee the approximated attention output equals to the attention computation on $\mathcal{I}_{t,\epsilon}$, RetrievalAttention uses a similar idea of FlashAttention [23] to combine $\mathbf{o}_{\mathcal{W}}$ and $\mathbf{o}_{\Omega}$ in the following equations:
@@ 7 $\Omega\xleftarrow{}$ VectorSearch($\mathbf{q}_{t}$);
Algorithm 1 summarizes the above design of RetrievalAttention and elaborate the procedure in an algorithm. At the beginning of each token generation, RetrievalAttention predicts active KV vectors and move them to GPU memory, and compute partial attention using the FlashAttention [23] kernel (1 - 1). In parallel with GPU computation, RetrievalAttention leverages the specially designed vector database to find the most relevant KV vectors to compute attention on CPU (1 - 1). Finally, RetrievalAttention combines the partial attention outputs on GPU and CPU using 4 and gets the approximated attention output (1).
@@ Appendix C Implementation
RetrievalAttention builds one individual vector index for the KV cache in one attention head. RetrievalAttention did several optimizations to optimize the prompt prefill, accelerate the vector search, and reduce the CPU memory usage.
@@
During the prefill phase, the full attention computation of the long context is required for generating the output vector for the next level of the LLM. Simultaneously, we move the KV vectors to the CPU side for the ANNS index building. To accelerate the overall prefill process, we overlap the cache movement to the CPU with the full attention computation on the GPU in a pipeline manner. To minimize peak GPU memory usage during the prefill phase, attention computation is performed sequentially across multiple attention heads. This approach only slightly impacts the attention computation speed, as longer prompts can fully leverage GPU parallelism with FlashAttention.
@@
To speed up the dynamic sparse attention computation on the CPU, we exploit the multi-thread parallelism in vector databases by leveraging the multi-core ability of modern CPU architecture. Specifically, since the computation of different attention heads is independent, we launch multiple threads for parallel searching across different vector indexes to reduce the overall latency on the CPU side. For grouped query attention (GQA) [36], although multiple query heads could share the same key-value vectors, we observe that the query vectors from different query heads in the same group exhibit different vector distributions. Therefore, we build one vector index for each query head to leverage the specific query distribution of each head.
@@
To reduce CPU memory consumption, the indexes in the same attention group share one copy of KV vectors by only storing the pointers to KV vectors in each index. In the future, we plan to utilize scalar quantization to further compress the KV vectors, implementing an 8-bit representation in place of the original FP32 format. This compression is promising to reduce memory usage while preserving computational efficiency. Importantly, our initial results demonstrate that this quantization approach does not compromise the accuracy of the model outcomes, maintaining performance equivalent to the full-precision representation.
@@ Appendix D Additional Related Work
Since the quadratic complexity of attention has become the bottleneck of LLM efficiency for long context applications, numerous works have studied to design sparse transformers to reduce the computational and memory complexity of the self-attention mechanism. Some works restrict the attention computation to predefined patterns including sliding windows [37], dilated windows [38], or mixture of different patterns [39, 40]. Some approaches use cluster-based sparsity based on hash value [41] or KNN algorithms [42, 43].These solutions either require pre-training a model from scratch or target limited scenarios like CPU-only, which do not work for our target to out-of-box usage of LLMs on the GPU-CPU architecture. Although some approaches [6, 7] exploit the dynamic sparse nature of LLMs, they often use some estimation using low-rank hidden states or post-statistical approaches, which incurs high overhead but with low accuracy. Moreover, all these approaches have to maintain full KV vectors on GPU with only accelerated inference by reduced memory movement, which does not solve the challenge of commodity GPUs with limited GPU memory.
Additionally, some approaches accelerate the inference by employing dynamically sparse attention patterns [22], separating the prefill and decoding stages [44, 20], and utilizing sequence parallelism [45, 46]. These methods are orthogonal to ours and can be in conjunction with our approach.

### v3

@@ A.4 Decoding Latency on A100
We test the generality of RetrievalAttention by measuring its performance on a server with one A100 (80GB) and one AMD EPYC 7V13 CPU with 24 cores and 220GB DRAM. We show the token-generation latency of different methods on three models in Table 8. Since the KV cache of full attention is disabled, all prompt tokens need to be recalculated during the decoding, incurring a very high decoding latency. By enabling the KV cache with the PageAttention optimization in vLLM, the decoding latency is significantly reduced. However, vLLM suffers from OOM issue with the increase of context length, which we elaborate further later. Other KV cache dropping or block retrieval methods including StreamingLLM, SnapKV, and InfLLM achieve faster decoding speed, but this is at the expense of a significant drop in model accuracy. In contrast, RetrievalAttention does not compromise generation accuracy while achieving much lower decoding latency than IVF and Flat because of the efficient mitigation of out-of-distribution problem.
@@ B.1 Formula of Combining Attention Results from the CPU and GPU Side
Attention operation is applied to the two sets of KV vectors separately on CPU and GPU, generating two partial attention outputs (denoted as $\mathbf{o}_{\mathcal{W}}$ and $\mathbf{o}_{\Omega}$, respectively). To guarantee the approximated attention output equals to the attention computation on $\mathcal{I}_{t,\epsilon}$, RetrievalAttention uses a similar idea of FlashAttention (Dao et al., 2022) to combine $\mathbf{o}_{\mathcal{W}}$ and $\mathbf{o}_{\Omega}$ in the following equations:
@@ 7 $\Omega\xleftarrow{}$ VectorSearch($\mathbf{q}_{t}$);
Algorithm 1 summarizes the above design of RetrievalAttention and elaborates the procedure in an algorithm. At the beginning of each token generation, RetrievalAttention predicts active KV vectors, moves them to GPU memory, and computes partial attention using the FlashAttention (Dao et al., 2022) kernel (#1 - #1). In parallel with GPU computation, RetrievalAttention leverages the specially designed vector database to find the most relevant KV vectors to compute attention on CPU (#1 - #1). Finally, RetrievalAttention combines the partial attention outputs on GPU and CPU using #4 and gets the approximated attention output (#1).
@@ Appendix C Implementation
RetrievalAttention builds one individual vector index for the KV cache in one attention head. RetrievalAttention has implemented several optimizations to optimize the prompt prefill, accelerate the vector search, and reduce CPU memory usage.
@@
During the prefill phase, full attention computation is required to generate the output vector for the next layer of the LLM. Simultaneously, we move the KV vectors to the CPU side for the ANNS index building. To accelerate the overall prefill process, we overlap the cache movement to the CPU with the full attention computation on the GPU in a pipeline manner. To minimize peak GPU memory usage during the prefill phase, attention computation is performed sequentially across multiple attention heads. This approach only slightly impacts the attention computation speed, as longer prompts can fully leverage GPU parallelism with FlashAttention.
@@
To speed up the dynamic sparse attention computation on the CPU, we exploit the multi-thread parallelism in vector databases by leveraging the multi-core ability of modern CPU architecture. Specifically, since the computation of different attention heads is independent, we launch multiple threads for parallel searching across different vector indexes to reduce the overall latency on the CPU side. For grouped query attention (GQA) (Ainslie et al., 2023), although multiple query heads could share the same key-value vectors, we observe that the query vectors from different query heads in the same group exhibit different vector distributions. Therefore, we build one vector index for each query head to leverage the specific query distribution of each head.
@@
To reduce CPU memory consumption, the indexes in the same attention group share one copy of KV vectors by only storing the pointers to KV vectors in each index. In the future, we plan to utilize scalar quantization to further compress the KV vectors, implementing an 8-bit quantization in place of the original FP16 format. This compression is promising to reduce memory usage while preserving computational efficiency. Importantly, our initial results demonstrate that this quantization approach does not compromise the inference accuracy, maintaining performance equivalent to the full-precision representation.
@@ Appendix D Additional Related Work
Since the quadratic complexity of attention has become the bottleneck of LLM efficiency for long context applications, numerous works have studied to design sparse transformers to reduce the computational and memory complexity of the self-attention mechanism. Some works restrict the attention computation to predefined patterns, including sliding windows (Child et al., 2019), dilated windows (Beltagy et al., 2020), or a mixture of different patterns (Zaheer et al., 2020; Ainslie et al., 2020). Some approaches use cluster-based sparsity based on hash value (Kitaev et al., 2020) or KNN algorithms (Bertsch et al., 2024; Mao et al., 2024). These solutions either require pre-training a model from scratch or target limited scenarios like CPU-only, which do not work for our target to out-of-box usage of LLMs on the GPU-CPU architecture. Although some approaches (Xiao et al., 2024a; Ribar et al., 2024) exploit the dynamic sparse nature of LLMs, they often use some estimation using low-rank hidden states or post-statistical approaches, which incurs high overhead but with low accuracy. Moreover, all these approaches have to maintain full KV vectors on GPU with only accelerated inference by reduced memory movement, which does not solve the challenge of commodity GPUs with limited GPU memory.
Additionally, some approaches accelerate the inference by employing dynamically sparse attention patterns (Jiang et al., 2024), separating the prefill and decoding stages (Zhong et al., 2024; Qin et al., 2024), and utilizing sequence parallelism (Jacobs et al., 2023; Liu et al., 2024a). These methods are orthogonal to ours and can be in conjunction with our approach.
Appendix E Additional Baselines
We compare RetrievalAttention with additional baselines InfiniGen and Quest on the RULER benchmark and show the results on Table 9. InfiniGen and Quest exhibit a noticeable drop in model accuracy compared to full attention. In contrast, RetrievalAttention performs best and achieves nearly the same accuracy as full attention across two benchmarks.
Table 9: Performance (%) of different methods in 128K context length on RULER.
Methods Act. Tokens S1 S2 S3 M1 M2 M3 MQ MV VT CW FW Q1 Q2 Avg.
Llama-3
FullAttention 128K 100.0 100.0 100.0 98.0 98.0 73.0 94.5 97.0 87.0 1.0 72.2 58.5 44.5 78.7
InfiniGen 2K 99.0 91.5 24.5 82.5 25.0 0.0 30.3 27.8 67.3 1.2 45.5 33.0 32.5 43.1
Quest 2K 100.0 100.0 98.5 98.5 36.5 0.0 48.9 64.3 89.4 1.0 64.5 45.0 39.5 60.5
Ours 640 + 100 100.0 100.0 100.0 99.0 98.0 45.0 92.8 93.0 88.0 1.1 49.3 60.5 44.5 74.7
Appendix F Dynamic Retrieval Budget Allocation
We investigated the impact of adjusting the retrieval budget according to the sparsity degree across layers, by adopting the budget allocation policy from PyramidIKV (Cai et al., 2024). Specifically, we compare the performance of the original RetrievalAttention with and without the PyramidKV-based budget allocation strategy on the InfiniteBench benchmark, as shown in RTable 2. Specifically, for the original RetrievalAttention, we set a fixed budget of 2000 tokens for all heads in all layers. In contrast, PyramidKV dynamically adjusts the retrieval size across different layers, allocating more in lower layers and less in higher ones.
The results in Table 10 shows that PyramidKV allocation strategy achieves better performance in Retr.KV tasks, though it slightly decreases performance in the En.QA task. On average, the accuracy slightly surpasses that of the original RetrievalAttention. This indicates that dynamic budget allocation is promising but may require task-specific allocation strategies.
Table 10: Performance (%) of RetrievalAttention and RetrievalAttention w/ PyramidKV in 128K context length.
Methods Retr.N Retr.P Retr.KV Code.D Math.F En.QA En.MC Avg.
Full Attention 100.0 100.0 17.5 19.0 39.5 9.1 68.0 50.4
RetrievalAttention 100.0 100.0 14.5 18.5 40.0 8.7 67.5 49.9
RetrievalAttention w/ PyramidKV
100.0 100.0 16.0 18.5 40.0 8.5 67.5 50.1
Appendix G Performance on the larger model
To demonstrate the generalizability of our methods on larger models, we evaluated our method on Llama-3-70B-262k using a server with eight 40GB A100 GPUs by partitioning the model by layers across GPUs. We choose the most complex task KV retrieval in $\infty$-Bench to stress test the efficiency of RetrievalAttention and other baselines.
[…]
Table 11: Performance (%) and decoding latency (s) in Llama-3-70B Model.
Full StreamingLLM Quest Flat RetrievalAttention
Accuracy 35.0 0.0 13.0 24.0 23.5
Decoding latency 248 0.14 1.36 5.68 1.62